# POPGym Arcade: Parallel Pixelated POMDPs

### v1

@@ Appendix E Environment Descriptions
    Battleship: The agent’s goal is to sink all the pre-placed ships to win the game. The agent controls a cursor that represents its current position on the grid. Each turn, the agent can either move the cursor one space or choose to "FIRE" the current grid cell. Each cell can be in one of two states: HIT or MISS. The agent receives a positive reward for hitting a ship, a neutral reward (0) for the first miss on an empty cell, and a negative reward for repeatedly firing a cell that has already been targeted.
@@
    Mine Sweeper: The agent’s goal is to hit all the non-mine cells to win the game, inspired by the classic computer game Minesweeper. Similar to Battleship environments, the agent controls a cursor to navigate the grid and can either move to an adjacent cell or "hit" the current cell each turn. Hitting a safe cell earns the agent a positive reward, while hitting a mine results in no reward (0) and ends the game. If the agent tries to hit a cell that has already been revealed, it receives a negative reward. This reward structure encourages the agent to explore efficiently, avoid mines, and minimize redundant actions to maximize its cumulative reward.
@@ Appendix G Network Architecture
Our Q-network combines spatial feature extraction with decision policy learning through a JAX/Equinox implementation. The network processes batches of 256×256 RGB images through four convolutional blocks, followed by three dense layers with intermediate normalization (Table 4).
@@
Conv2D-1 Channels: $3,64$, Kernel: $7\times 7$, Stride: 2 LeakyReLU
@@
Conv2D-4 Channels: $256,512$, Kernel: $3\times 3$, Stride: 2 LeakyReLU
@@
Conv2D-1 Channels: $3,64$, Kernel: $7\times 7$, Stride: 2 LeakyReLU
@@
Conv2D-4 Channels: $256,512$, Kernel: $3\times 3$, Stride: 2 LeakyReLU
RNN Cell Input, Hidden, Output: $517,512,256$, Num Layer: 1
Dense-1 Features: $256,256$ LeakyReLU
LayerNorm Shape: 256
Dense-2 Features: $256,5$
Appendix H Experiment Hyperparameters
@@
TOTAL_TIMESTEPS 10e6,20e6,100e6
TOTAL_TIMESTEPS_DECAY 1e6,2e6,20e6

### v5

@@ Appendix E Environment Descriptions
    Battleship: The agent’s goal is to sink all the pre-placed ships to win the game. The agent controls a cursor that represents its current position on the grid. Each turn, the agent can either move the cursor one space or choose to ”FIRE” the current grid cell. Each cell can be in one of two states: HIT or MISS. The agent receives a positive reward for hitting a ship, a neutral reward (0) for the first miss on an empty cell, and a negative reward for repeatedly firing a cell that has already been targeted.
@@
    Mine Sweeper: The agent’s goal is to hit all the non-mine cells to win the game, inspired by the classic computer game Minesweeper. Similar to Battleship environments, the agent controls a cursor to navigate the grid and can either move to an adjacent cell or ”hit” the current cell each turn. Hitting a safe cell earns the agent a positive reward, while hitting a mine results in no reward (0) and ends the game. If the agent tries to hit a cell that has already been revealed, it receives a negative reward. This reward structure encourages the agent to explore efficiently, avoid mines, and minimize redundant actions to maximize its cumulative reward.
@@ Appendix G Network Architecture
Our Q-network combines spatial feature extraction with decision policy learning through a JAX/Equinox implementation. The network processes batches of $128\times 128$ RGB images through four convolutional blocks, followed by three dense layers with intermediate normalization (Table 6).
@@
Conv2D-1 Channels: $3,64$, Kernel: $5\times 5$, Stride: 2 LeakyReLU
@@
Conv2D-4 Channels: $256,512$, Kernel: $1\times 1$, Stride: 1 LeakyReLU
@@
Conv2D-1 Channels: $3,64$, Kernel: $5\times 5$, Stride: 2 LeakyReLU
@@
Conv2D-4 Channels: $256,512$, Kernel: $1\times 1$, Stride: 1 LeakyReLU
RNN Cell Input, Hidden, Output: $517,512,256$, Num Layer: 2
Dense Features: $256,5$
Appendix H Recurrent Models
In this section, we describe all recurrent models used in our work.
Linear Transformers
Standard transformers have quadratic space complexity from the outer product of keys and queries. The Fast Autoregressive Transformer (FART) (Katharopoulos et al., 2020) replaces softmax attention with a kernelized attention mechanism to achieve linear space complexity and logarithmic time complexity via associative recurrent updates
$h_{t}=h_{t-1}+\phi(W_{k}x_{t})\phi(W_{v}x_{t})^{\top}$ $\hat{s}_{t}=\mathrm{MLP}\left(x+\frac{\phi(W_{q}x_{t})^{\top}h_{t}}{\phi(W_{q}x_{t})\cdot\sum_{i=0}^{t}\phi(W_{k}x_{i})}\right).$ (10)
Here, $\phi(x)=1+\mathrm{ELU}(x)$ represents a kernel-space projection and the recurrent state $h$ represents the attention matrix. $W_{k},W_{v},W_{q}$ represent the key, query, and value projection parameters. To compute $\hat{s}_{t}$, we multiply the recurrent attention matrix by the query vector and normalize by a scalar.
State-Space Models
$h_{t}=\overline{W}_{A}h_{t-1}+\overline{W}_{B}x_{t}$ $\hat{s}_{t}=\overline{W}_{C}h_{t}+\overline{W}_{D}x_{t},$ (11)
where $\overline{W}_{A},\overline{W}_{B},\overline{W}_{C},\overline{W}_{D}$ are discretizations of carefully initialized trainable parameters $W_{A},W_{B},W_{C},W_{D}$. In practice, we initialize $W_{A},W_{B},W_{C},W_{D}$ deterministically. Deterministic initialization applied to many consecutive SSM layers can result in instabilities, and so Orvieto et al. (2023) proposes a meticulously derived random initialization and new parameterization of SSM parameters. They call their method the Linear Recurrent Unit (LRU).
Minimal Recurrent Networks
$h_{t}=(1-\sigma(W_{1}x_{t}+b_{1}))\odot h_{t-1}+\sigma(W_{1}x_{t}+b_{1})\odot\tanh(W_{2}x_{t}+b_{2})$ $\hat{s}_{t}=h_{t},$ (12)
with trainable parameters $W_{1},b_{1},W_{2},b_{2}$, sigmoid function $\sigma$, and elementwise product $\odot$. The authors find that the MinGRU outperforms multiple SSM variants across offline reinforcement learning tasks, using an offline decision transformer (Chen et al., 2021) framework.
Appendix I Experiment Hyperparameters
Experiment hyperparameters were taken from the original PQN paper (Gallici et al., 2024). We slightly modified the hyperparameters to minimize training time and produce more stable return curves.
@@ Appendix I Experiment Hyperparameters
TOTAL_TIMESTEPS 10e6,20e6
TOTAL_TIMESTEPS_DECAY 1e6,2e6
@@
Appendix J Compute Resources
We used approximately 98 days equivalent training time on an RTX4090 GPU to produce a little over 1,000 total runs. We used an unknown amount of additional resources to tune hyperparameters and rerun experiments after fixing bugs.